whats molarity and the factors affecting it?
does temperature affect it.

Grade:10

2 Answers

utkarsh
7 years ago
Molarity = net number of moles of solute/ volume of solution (in litres) .. Let x be Number of moles of HCl (in mixture)= no of moles of HCl from solution A+solution B. Then x = (0.5* 0.750)+(2*0.250)=0.875. And let y= total volume of mixture = 1 litre. Therefore molarity =x/y = 0.875/1= 0.875. #always take volume in litres.
Umakant biswal
7 years ago
Molarity is a unit of concentration measuring the number of moles of a solute per liter of solution..
it can be calculated by the following formula 
no of moles of solute / no of litres of solution 

Molarity is affected by temperature because it is based on the volume of the solution, and the volume of a substance will be affected by changes in temperature..
for that reason molality is favoured over molarity to calculate the concentration terms 
HOPE IT CLEARS YOUR DOUBT